Find equation of line passing through the point (5,-2) and (-3,4)
I got the answer y=(-3/4)x+7/4 but I don't know how
Please show me step by step to get the answer. thanks

Answer by nerdybill(7384)

Find equation of line passing through the point (5,-2) and (-3,4)
.
First, find the slope:
m = (y2-y1)/(x2-x1)
m = (4-(-2))/(-3-5)
m = (4+2)/(-3-5)
m = (6)/(-8)
m = -3/4
.
Next, use one of the two given points (5, -2) and the slope above -- plug it into the "point-slope" form:
y - y1 = m(x - x1)
y - (-2) = (-3/4)(x - 5)
y + 2 = (-3/4)x + 15/4
y = (-3/4)x + 15/4 - 2
y = (-3/4)x + 15/4 - 8/4
y = (-3/4)x + 7/4
.
So, YES, you are correct!